it takes John 7 hours to complete a journey of 640 km. He travels by train for the first 4 hours of the journey and the last part by car. the car travels 50 km/h faster than the train. What is the speed of the train
Answer by ikleyn(52754)

.
Let x be the speed of the train, in km/h.


Then the speed by the car is (x+50) km/h.


The full distance equation is


    4x + 3*(x+50) = 640       <<<---=== notice that the time to travel by the car is  7-4 = 3 hours


    4x + 3x + 150 = 640


    7x = 640-150 = 490  ====>  x = 490/7 = 70.


Answer.   The speed of the train is  70 km/h.

Solved.